OpenStudy (anonymous):

Need help; Graph the system of inequalities x + 4y >-8; y<3

OpenStudy (anonymous):

solve for y 4y > -8-x y > -2-x/4 now draw two graphs y = 3 (horizintal line) y = -2-x/4 (line with negative slope, not steep) Look at the region below y = 3 and above y = -2-x/4, if those regions intersect, shade that and that should be your answer :)